OpenStudy (anonymous):

Find the simple interest earned after 12 years on an investment of $5,000 at 4% interest earned annually.

OpenStudy (anonymous):

can someone help me with this plz

OpenStudy (anonymous):

Formula for simple interest: A = P(1 + rt), so A = 5000(1 + (0.04 × 12)) = 7400 A = $ 7,400.00